# Refinement Module based on Parse Graph of Feature Map for Human Pose Estimation

### v1

@@ IV Experiments
TABLE V: Ablation experiment comparison on the MPII test set. Set up unsupervised RMPG modules only. The supervised RMPG module settings remain unchanged, setting $gp=[5,2]$. Please note that multi-scale testing [58] is not used in ablation experiments. $\dagger$ denotes the results of our reimplementation.
@@
Figure 4: (a) An input image of our framework. (b) Visualization of context relations among all joints in (a). Each row represents the relation between a corresponding joint and all joints. The higher the score, the stronger the relation. From the visualization of the context relations in joints, we notice that the joint tends to have a higher score with itself when it is visible, such as left hip, left wrist, and left ankle. On the contrary, the joints tend to have a higher score with other joints when they are occluded or invisible, such as right wrist, right ankle, and right shoulder.
@@ IV-B Implementation details
We use $gp$ to represent the settings of the tree structure in RMPG, including the number of nodes and depth. For example, $gp=[n_{1},n_{2},n_{3}]$ represents a tree with a depth of $3$, in which there are $n_{1}$ nodes in the first level, each node in the first level has $n_{2}$ child nodes, and each node in the second level has $n_{3}$ child nodes. To be specific, Fig. 2b is $gp=[3]$ and Fig. 2c is $gp=[3,3]$. In our framework, for supervised RMPG modules, setting $gp=[5,2]$, and for unsupervised RMPG modules, setting $gp=[2,2]$.
@@ IV-C Benchmark results
Our framework is divided into small model and large model. The main difference between them is that the number of convolution layers are different, resulting in changes in complexity, but there is no difference in the overall structure. Their parameter quantities are shown in Table I. The parameter size of our small model is $37.1M$, which is $8.6M$ more than the baseline HRNet-W32. The parameter size of our large model is $50.7M$, which is $22.2M$ more than the benchmark model HRNet-W32, $30.3M$ less than the method of [2], and $35.3M$ less than VITPose-B [23].
COCO keypoint detection benchmark. Tables I and II show the results of our method and existing advanced methods on the coco validation and test-dev sets. For the COCO validation set, Our small model achieve $75.3$ MAP with the input size of $256\times 192$ and $76.9$ MAP with the input size of $384\times 288$, which are $0.9$ and $1.1$ higher than the baseline HRNet-W32 respectively. Our large model achieve $75.8$ MAP with the input size of $256\times 192$ and $77.4$ MAP with the input size of $384\times 288$, which are $1.4$ and $1.6$ higher than the baseline HRNet-W32 respectively, and $0.2$ and $0.5$ higher than the method of PGBS [2] respectively. Compared with VITPose-B [23], our large model, with low parameters, achieves the same accuracy as VITPose-B. For the COCO test-dev set, our small model achieve $74.4$ MAP with the input size of $256\times 192$ and $75.8$ MAP with the input size of $384\times 288$, which are $1.1$ and $0.9$ higher than the baseline HRNet-W32 respectively. Furthermore, our small model exceeds PGBS [2] by 0.1 when the input is $384\times 288$. Our large model achieves $75.0$ MAP with the input size of $256\times 192$ and $76.3$ MAP with the input size of $384\times 288$, which are $1.5$ and $1.4$ higher than the baseline HRNet-W32 respectively, and $0.4$ and $0.6$ higher than the method of PGBS [2] respectively.
@@
MPII benchmark. As shown in Table IV, the PCKh@0.5 results of our method and other state-of-the-art methods on the MPII test set. By reproducing HRNet [31] with the input size of $256\times 256$, we achieve a $91.5$ PCKh@0.5 score with multi-scale testing [58]. Our small model achieves a $92.1$ PKCh@0.5 score with the input size of $256\times 256$, which is $0.6$ higher than the baseline HRNet. Our big model achieves a $92.3$ PKCh@0.5 score with the input size of $256\times 256$, which is $0.8$ higher than the baseline HRNet and $0.2$ higher than the method of [2].
@@ IV-D Ablation study
Our ablation experiments are made on the MPII test set without multi-scale testing [58]. Mean PCKh@0.5 is used as the evaluation metric. In ablation experiments, our small model is used. All results are obtained with the input size of $256\times 256$. Table V shows the PCKh@0.5 results of our ablative experiment. We reimplement the HRNet-W32 with the input size $256\times 256$ and it achieves a $91.2$ PCKh@0.5 score without multi-scale testing [58]. In addition, we obtain a result of $91.3$ PCKh@0.5 score without multi-scale testing using the trained weights and code provided by HRNet [31].
It can be seen from Table V that an appropriate increase in the number of leaves is beneficial for improving the results, such as the comparison between $gp=[1]$ and $gp=[4]$. An appropriate increase in depth is also beneficial for improving the results, such as $gp=[2,4]$ and $gp=[2,2]$. In addition, we also explore more about the impact of the number and depth of branches on the results in the next subsection.
@@ IV-E RMPG performance on different methods
The result of the RMPG on different methods as shown in Table VI, for the method of Hourglass, SimpleBaselines and ViTPose, the results are improved after adding the RMPG. To be specific, for the method of SimpleBaselines, we achieve 0.7 MAP improvement after introducing the RMPG module ($gp=[2,2]$ and $gp=[4,4]$) and 0.8 MAP improvement after introducing the RMPG module ($gp=[4,2]$) when the backbone is ResNet-50. What’s more, we achieve 0.3 MAP improvement after introducing the RMPG module ($gp=[2,2]$) and 0.6 MAP improvement after introducing the RMPG module ($gp=[4,2]$) when the backbone is ResNet-101. For the method of Hourglass [22], we achieve 1.4 MAP improvement after introducing the RMPG module ($gp=[2,2]$) and 1.2 MAP improvement after introducing the RMPG module ($gp=[4,4]$, $gp=[4,2]$ and $gp=[2,2,2]$) when the backbone is Hourglass-52. For the method of ViTPose [23], we achieve 0.3 MAP improvement after introducing the RMPG module ($gp=[2,2]$) when the backbone is ViTPose-B. The results of other $gp$ settings are shown in the Table VI.
The above results prove the effectiveness of the RMPG module. The parameters of the RMPG will increase as the channels of the feature map increases. For example, for ViTPose-B, the number of channels increases significantly compared to other methods in Table VI, from 256 to 768. Therefore, after the introduction of the RMPG module, its parameters increased significantly.
@@ V Conclusion
We designed the RMPG module based on parse graphs, which can obtain context relations between sub-feature maps, thereby optimizing the network for better performance. In addition, RMPG can help to explicitly model context relations and hierarchies in parse graph of body structure. The experimental results demonstrate the effectiveness of our method, and we hope that the RMPG module can be widely used in various vision tasks.

### v3

@@ IV Experiments
TABLE V: Ablation experiment comparison on the MPII test set. Set up unsupervised RMPG modules only. The supervised RMPG module settings remain unchanged, setting $gp=[5,2]$. $\dagger$ denotes the results of our reimplementation. $CR$ means context relations. Please note that multi-scale testing [52] is not used in ablation experiments.
@@ IV-B Implementation details
We use $gp$ to represent the settings of the tree structure in RMPG, including the number of nodes and depth. For example, $gp=[n_{1},n_{2},n_{3}]$ represents a tree with a depth of $gp$ list length $3$, in which there are $n_{1}$ nodes in the first level, each node in the first level has $n_{2}$ child nodes, and each node in the second level has $n_{3}$ child nodes. To be specific, Fig. 2b is $gp=[3]$ and Fig. 2c is $gp=[3,3]$. In our network, for supervised RMPG modules, setting $gp=[5,2]$, where 5 is the number of sub-feature maps after the first decomposition corresponding to the 5 body parts, and they are supervised by different body parts during training. For unsupervised RMPG modules, setting $gp=[2,2]$.
@@ IV-C Benchmark results
Our network is divided into small model and large model. The main difference between them is that the number of convolution layers are different, resulting in changes in complexity, but there is no difference in the overall structure
COCO keypoint detection benchmark. Tables I and II show the results of our method and existing advanced methods on the coco validation and test-dev sets. For the COCO validation set, Our small model achieve $75.3$ MAP with the input size of $256\times 192$ and $76.9$ MAP with the input size of $384\times 288$, which are $0.9$ and $1.1$ higher than the baseline HRNet-W32 respectively. Our large model achieve $75.8$ MAP with the input size of $256\times 192$ and $77.4$ MAP with the input size of $384\times 288$, which are $1.4$ and $1.6$ higher than the baseline HRNet-W32 respectively, and $0.2$ and $0.5$ higher than the method of PGBS [1] respectively. For the COCO test-dev set, our small model achieve $74.4$ MAP with the input size of $256\times 192$ and $75.8$ MAP with the input size of $384\times 288$, which are $1.1$ and $0.9$ higher than the baseline HRNet-W32 respectively. Furthermore, our small model exceeds PGBS [1] by 0.1 when the input is $384\times 288$. Our large model achieves $75.0$ MAP with the input size of $256\times 192$ and $76.3$ MAP with the input size of $384\times 288$, which are $1.5$ and $1.4$ higher than the baseline HRNet-W32 respectively, and $0.4$ and $0.6$ higher than the method of PGBS [1] respectively. Compared with PGBS and ViT-B, our network can achieve better results with lower complexity.
@@
MPII benchmark. As shown in Tabel IV, the PCKh@0.5 results of our method and other state-of-the-art methods on the MPII test set. By reproducing HRNet [20] with the input size of $256\times 256$, we achieve a $91.5$ PCKh@0.5 score with multi-scale testing [52]. Our small model achieves a $92.1$ PKCh@0.5 score with the input size of $256\times 256$, which is $0.6$ higher than the baseline HRNet and $1.0$ higher than TokenPose [21]. Our big model achieves a $92.3$ PKCh@0.5 score with the input size of $256\times 256$, which is $0.8$ higher than the baseline HRNet and $0.2$ higher than the method of [1].
@@ IV-D Ablation study
Our network ablation experiments are made on the MPII test set without multi-scale testing [52]. Mean PCKh@0.5 is used as the evaluation metric. In ablation experiments, our small model is used. All results are obtained with the input size of $256\times 256$. Tabel V shows the PCKh@0.5 results of our ablative experiment.
It can be seen from Tabel V that an appropriate increase in the number of leaves is beneficial for improving the results, such as the comparison between $gp=[1]$ and $gp=[4]$. An appropriate increase in depth is also beneficial for improving the results, such as $gp=[2,4]$ and $gp=[2,2]$. In addition, we also explore more about the impact of the number and depth of branches on the results in the next subsection.
In order to verify the effectiveness of the context relations, we also remove the context relations in the parse graph. When $gp=[2,2]$ without context relations, we get 91.4 PCKh@0.5, which is reduced by 0.5. This proves that the context relations is valid.
@@ IV-E RMPG performance on different methods
The result of the RMPG on different methods as shown in Tabel VI, for the method of Hourglass, SimpleBaselines and ViTPose, the results are improved after adding the RMPG. To be specific, for the method of SimpleBaselines, we achieve 0.7 MAP improvement after introducing the RMPG module ($gp=[2,2]$ and $gp=[4,4]$) and 0.8 MAP improvement after introducing the RMPG module ($gp=[4,2]$) when the backbone is ResNet-50. What’s more, we achieve 0.3 MAP improvement after introducing the RMPG module ($gp=[2,2]$) and 0.6 MAP improvement after introducing the RMPG module ($gp=[4,2]$) when the backbone is ResNet-101. For the method of Hourglass [18], we achieve 1.4 MAP improvement after introducing the RMPG module ($gp=[2,2]$) and 1.2 MAP improvement after introducing the RMPG module ($gp=[4,4]$, $gp=[4,2]$ and $gp=[2,2,2]$) when the backbone is Hourglass-52. For the method of ViTPose [19], we achieve 0.3 MAP improvement after introducing the RMPG module ($gp=[2,2]$) when the backbone is ViTPose-B. The results of other $gp$ settings are shown in the Tabel VI. The above results prove the effectiveness of the RMPG module.
@@ V Conclusion
The RMPG module based on parse graphs can obtain context relations between sub-feature maps, optimizing the network for better performance. In addition, RMPG can help to explicitly model context relations and hierarchies with lower parameters in parse graph of body structure. The experimental results demonstrate the effectiveness of our method, and we hope that the RMPG module can be widely used in various vision tasks.